QUESTION 3

  1. A colleague of yours has a K100,000-00, 2 years treasury Bond maturing in 12 months, issued at a fixed coupon of 10%, payable annually. He informs you that he has an urgent need of money and wants to sell you the Bond.

  1. What’s the maximum price you would offer assuming the yield on a 12 months treasury bill is currently at 12%?

                                                                        [04 Marks]

  1. Briefly discuss how you may be affected by inflation over the holding period to maturity.

                                                                        [06 Marks]

  1. A 273 Days Treasury bill of K1,000 Face Value is currently on offer at K840.33.

  1. Calculate its Yield to maturity (YTM) at this price.

                                                                        [02 Marks]

  1. Calculate its Effective Annual Return (APR).

                                                                        [03 Marks]

  1. Briefly explain why Treasury bills are treated as risk-free securities.                                                                                          [05 Marks]

Total 20 Marks
Answer #1

Answer a1) Price of 1 year Bond = Coupon + maturity (1+Y) = (1000+100000)/(1+0.10) = K 91818.18

Answer a2)

The twin factors that influence a security's value are expansion and changing loan fees. An ascent in either financing costs or the inflation rate will in general reason security costs to drop. Swelling and loan costs carry on comparably to security yields, moving the other way from security costs.

inflation rate disintegrates the buying intensity of a security's future incomes. Put basically, the higher the present pace of swelling and the higher the (normal) future paces of expansion, the higher the yields will ascend over the yield bend, as financial specialists will request this better return to make up for swelling hazard.

Answer b1)

1000 - Price yield = 1 Price - * 100

Yield = (1000-840.33)/840.33 * 100= 19.001%

Answer b2) Annualized yield =365 yield * - maturity day

Annualized yield = 19.001%*365/273 = 25.404%

Answer b3) A risk free resource has a specific future return. Treasuries (particularly T-bills) are viewed as free from any type of risk in light of the fact that the U.S. government backs them. Since they are so sheltered, the arrival on risk free resources is exceptionally near the present financing cost.